### v1

# Conformal Data for the O(3) Wilson-Fisher CFT from Fuzzy Sphere Realization of Quantum Rotor Model

Conformal Data for the O(3) Wilson-Fisher CFT from Fuzzy Sphere Realization of Quantum Rotor Model
Arjun Dey Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Loic Herviou Univ. Grenoble Alpes, CNRS, LPMMC, 38000 Grenoble, France    Christopher Mudry Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Andreas Martin Läuchli Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland
Аннотация
We present a model for strongly interacting fermions with internal O(3) symmetry on the fuzzy-sphere that (i) preserves the rotational symmetry of the fuzzy sphere and (ii) undergoes a quantum phase transition in the (2+1)-dimensional O(3) Wilson–Fisher universality class. Using exact diagonalization (ED) and density matrix renormalization group (DMRG), we locate the quantum critical point via conformal perturbation theory and obtain scaling dimensions from finite-size spectra. We identify 24 primary operators and determine some of their operator product expansion coefficients through first-order conformal perturbation theory. The results are benchmarked against conformal bootstrap and large quantum-number expansions and reveal a weakly irrelevant operator that plays a role in dimerized antiferromagnets. Our work establishes the fuzzy sphere as a general framework for quantitatively accessing conformal data in non-Abelian conformal field theories (CFTs).
Introduction.— CFTs in $(d+1)$-dimensional spacetime [(d+1)(d+1)D] are fully specified by their conformal data, scaling dimensions, and operator product expansion (OPE) coefficients, which govern universal critical behavior, and some measurable responses Polyakov (1970); Di Francesco et al. (1997); Rychkov (2017); Cardy (1996); Goldenfeld (2018); Witczak-Krempa et al. (2012); Sachdev (2011). While (1+1)D CFTs Di Francesco et al. (1997); Belavin et al. (1984) enjoy remarkable theoretical completeness through Virasoro algebra Virasoro (1970) and integrability Takahashi (1999), the landscape of (2+1)D CFTs remains largely unexplored, presenting rich opportunities for discovering new universal physics and testing fundamental principles of quantum field theory.
Standard approaches have limitations when it comes to accessing detailed information about operator spectra and OPE data of (2+1)D CFTs. $(i)$ Large-scale Monte Carlo simulations on lattices are effective for extracting critical exponents Hasenbusch (2022), but they face significant challenges when trying to resolve additional low-lying operators or obtain OPE coefficients. $(ii)$ The conformal bootstrap has achieved remarkable precision for the Ising universality class El-Showk et al. (2012); Kos et al. (2016), providing tight bounds and detailed information for the lowest-lying operators. However, for theories with non-Abelian symmetries such as O(3), it typically yields constraints for only a limited set of low-lying operators, especially in higher-spin sectors, and becomes increasingly computationally intensive. Moreover, it does not offer a direct Hamiltonian realization. $(iii)$ Analytic approaches often produce asymptotic series that require resummation and extrapolation to the physical dimension of space or the desired symmetry group, with growing uncertainties for spinning/higher-rank operators and specific OPE coefficients Derkachov and Manashov (1997); Lang and Rühl (1993). $(iv)$ Fuzzy sphere regularization Zhu et al. (2023); Hu et al. (2023); Han et al. (2023); Hu et al. (2024); Zhou et al. (2024a); Hu et al. (2025); Cuomo et al. (2024); Zhou and Zou (2025); Fardelli et al. (2025); Zhou et al. (2024b); He (2025); Taylor et al. (2025) provides a direct Hamiltonian approach to studying (2+1)D CFTs through diagonalization of quantum systems on a sphere, offering access to both spectra and correlation functions while preserving full symmetries. A prior fuzzy-sphere study of the (2+1)D CFT with internal O(3) symmetry confirmed feasibility but was limited by Hilbert-space growth and identified only a limited number of operators Han et al. (2024).
Here, we introduce a fuzzy-sphere realization (FZR) of a truncated quantum rotor model (TQRM) Sachdev (2011); Zhu et al. (2023) that retains the full internal O(3) and spatial SO(3) symmetry of the (2+1)-dimensional O(3) Wilson–Fisher universality class and is tractable to ED and DMRG. Building on the well‐known correspondence between O(3) quantum rotors and O(3) quantum criticality Sachdev (2011), the idea is to use a conformal map to identify a compactification of two-dimensional space with the unit 2-sphere $\mathbb{S}^{2}$ and to leverage the state-operator correspondence to extract conformal data. This is done by relating the finite-size gaps of the FZR of the TQRM to the scaling dimensions of the CFT operators. We identify 24 low-lying primary operators, with their scaling dimensions summarized in Tables 1 and 2 (Appendix A). Selected OPE coefficients are also determined. The full set of scaling dimensions and OPE values can be found in the Supplementary Material Sup . We benchmark our results with bootstrap results Chester et al. (2021) and large-charge predictions Cuomo (2020); Cuomo et al. (2021), wherever they are available.
Quantum rotor model.— A rigid rotor, a particle of mass $\mu$ constrained to move on the surface of a sphere of radius $R$, serves as the prototypical quantum model of rotational motion. Its energy is purely kinetic, arising from quantized angular momentum with moment of inertia $I=\mu\,R^{2}>0$. On a square lattice, an array of such rotors with nearest-neighbor ferromagnetic Heisenberg coupling $J>0$ realizes an O(3) magnetic transition when the dimensionless coupling IJI\,J is tuned to quantum criticality Sachdev (2011).
The full two-dimensional interacting lattice Hamiltonian for O(3) quantum rotors reads $\widehat{H}_{\infty}:=\widehat{H}_{\mathrm{kin}}+\widehat{H}_{\mathrm{int}},$
$\begin{split}&\widehat{H}_{\mathrm{kin}}:=\frac{1}{2I}\sum_{i=1}^{N}\widehat{\mathbf{L}}_{i}^{2},\qquad\widehat{H}_{\mathrm{int}}:=-J\sum_{\langle ij\rangle}\hat{\mathbf{n}}_{i}\cdot\hat{\mathbf{n}}_{j},\end{split}$ (1a)
where the pair of operator-valued 3-component vectors $(\widehat{\mathbf{L}}_{i},\hat{\mathbf{n}}_{i})$ obey the equal-time O(3) rotor algebra
$\begin{split}&[\widehat{L}_{i}^{a},\widehat{L}_{j}^{b}]:=\mathrm{i}\hbar\,\epsilon^{abc}\,\widehat{L}_{i}^{c}\,\delta_{ij},\\ &[\widehat{L}_{i}^{a},\hat{n}_{j}^{b}]:=\mathrm{i}\hbar\,\epsilon^{abc}\,\hat{n}_{i}^{c}\,\delta_{ij},\quad a,b=1,2,3\equiv x,y,z,\\ &[\hat{n}_{i}^{a},\hat{n}_{j}^{b}]:=0,\end{split}$ (1b)
with the local constraints
$\widehat{\mathbf{L}}_{i}^{2}=\hbar^{2}\,l(l+1)\,\widehat{\mathbb{I}},\qquad\hat{\mathbf{n}}_{i}^{2}:=\widehat{\mathbb{I}}$. (1c)
Here, $l\in\mathbb{Z}_{\geq 0}$ fixes a global representation of each quantum rotor, $\langle ij\rangle$ are nearest-neighbors sites on the square lattice with $N$ the number of sites.
In the limit $I\,J=\infty$, $\widehat{H}_{\infty}$ simplifies to the classical O(3) ferromagnetic Heisenberg model
$\widehat{H}_{\mathrm{int}}=-J\sum_{\langle ij\rangle}\mathbf{n}_{i}\cdot\mathbf{n}_{j},\qquad\mathbf{n}_{i}^{2}=1,$ (2)
with the ferromagnetic ground state $|\mathbf{n}_{1},\ldots,\mathbf{n}_{N}\rangle=|\mathbf{n},\ldots,\mathbf{n}\rangle$ whereby $\hat{\mathbf{n}}_{i}\,|\mathbf{n}_{1},\ldots,\mathbf{n}_{N}\rangle=\mathbf{n}_{i}\,|\mathbf{n}_{1},\ldots,\mathbf{n}_{N}\rangle$. In the limit IJ=0I\,J=0, $\widehat{H}_{\infty}$ simplifies to
$|\displaystyle\widehat{H}_{\mathrm{kin}}=\sum_{i=1}^{N}\sum_{l=0}^{\infty}\sum_{m_{i}=-l}^{+l}\frac{\hbar^{2}\,l(l+1)}{2I}\ket{l,m_{i}}\bra{l,m_{i}}$ (3)
[…]
Fuzzy-sphere many-body electronic Hamiltonian.— To access conformal data, we map the quantum rotor model to a system of $N$ fermions with internal degrees of freedom, moving on a sphere Haldane (1983) and projected to the lowest Landau level (LLL). In this limit, the noncommuting coordinate operators generate the fuzzy-sphere algebra Madore (1992), thereby providing a natural short-distance cutoff (the magnetic length) that preserves rotational symmetry.
We focus on the case where each fermion has $F=4$ flavors and the system reside at quarter filling. At quarter filling, strong interactions enforce (on average) single occupancy per LLL orbital, analogous to the large-$U$ limit of the Hubbard model.
The microscopic fuzzy-sphere Hamiltonian that we shall study numerically is defined to be
$\widehat{H}_{\mathrm{fzs}}:=\widehat{P}_{\mathrm{LLL}}\left(u\,\widehat{H}_{\mathrm{Hub}}-v\,\widehat{H}_{\mathrm{Heis}}+h\,\widehat{H}_{\mathrm{trans}}\right)\widehat{P}_{\mathrm{LLL}}.$ (4)
[…]
Таблица 1: Summary of selected primary operators. $S$ labels the internal SO(3) spin representation, $L$ the SO(3) angular momentum on the fuzzy sphere, and $\pm$ the $\mathbb{Z}_{2}$ parity (improper part) of the internal O(3) symmetry, so that states are labeled by $S^{\pm}$ and $L$. The index I orders energy eigenstates within each symmetry sector. Scaling dimensions are extracted via ED and DMRG, and are compared to conformal bootstrap (CB); asterisks denote exact values, and parentheses indicate system size. The two bold-faced entries in the $\Delta$ (CB) column are used as input.
$S^{\pm}$ $L$ II $\mathrm{o}$ $\Delta$(ED) $\Delta$(DMRG) $\Delta$(CB)
$1^{-}$ 0 1 $\sigma$ 0.51893 (12) 0.51893 (28) 0.518936 Chester et al. (2021)
$0^{+}$ 0 2 $\varepsilon$ 1.56189 (12) 1.61781 (26) 1.59488 Chester et al. (2021)
$0^{+}$ 0 4 $\varepsilon^{\prime}$ 3.77279 (12) 3.80177 (14) 3.76680 Henriksson (2023)
$0^{+}$ 2 1 $T_{\mu\nu}$ 2.97749 (12) 3.01062 (20) $3^{*}$
$^{+} 1 1$ $j_{\mu}$ 1.88253 (12) 1.97204 (26) $2^{*}$
$2^{+}$ 0 1 $t_{(2)}$ 1.25407 (12) 1.23767 (26) 1.20954 Chester et al. (2021)
$4^{+}$ 0 1 $t_{(4)}$ 3.26858 (11) 3.16459 (26) 2.99056 Chester et al. (2021)
Рис. 1: Panel (a) shows the dependence on $h$ of the relevant coupling $g_{\varepsilon}(h)$ for a given $R\propto\sqrt{N}$. The critical point $h_{\mathrm{c}}$ is the root to $g_{\varepsilon}(h)=0$ after the thermodynamic limit $N\uparrow\infty$ has been taken. Panel (b) shows the dependence of the root to $g_{\varepsilon}(h)=0$ as a function of $1/\sqrt{N}\propto 1/R$. Fitting this dependence with an appropriate power law and using the first irrelevant anomalous dimension $0.7688$ Henriksson (2023), we extract the critical value $h_{\mathrm{c}}=14.992\pm 0.018$ and the critical exponent $1/\nu=1.3908\pm 0.0852$, which is in close agreement with the known value $1.406$ Campostrini et al. (2002).
Symmetries.— Hamiltonian (4) exhibits three symmetries: (a) spatial rotational symmetry SO(3), arising from conservation of angular momentum with the quantum number $L$; (b) internal SO(3) spin symmetry with the quantum number $S$; (c) and a $\mathbb{Z}_{2}$ symmetry corresponding to the conservation of the parity of the number of electrons occupying $p$ orbitals in internal spin space. The quantum number for the $\mathrm{SO}(3)\times\mathbb{Z}_{2}\cong\mathrm{O}(3)$ is denoted by $S^{\pm}$, where the superscript denotes the $\mathbb{Z}_{2}$ parity.
Accessing quantum criticality through the state-operator correspondence.— We first assume the existence of a quantum critical point $\delta u=\delta v=\delta h=0$ at which $\widehat{H}_{\mathrm{fzs}}(R)$ exhibits the conformal invariance of the (2+1)D O(3) Wilson-Fisher universality class in the thermodynamic limit $R\propto\sqrt{N}\uparrow\infty$. If we denote eigenstates of $\widehat{H}_{\mathrm{fzs}}(R)$ by the label $\mathrm{o}$ (see Table 1), the state-operator correspondence Rychkov (2017) implies that the eigenenergy $E_{\mathrm{o}}(R)$ measured relative to that of the ground state $E_{0}(R)$ scales like
$\delta E_{\mathrm{o}}(R)\equiv E_{\mathrm{o}}(R)-E_{0}(R)\sim\frac{c}{R}\,\Delta_{\mathrm{o}}\qquad(\hbar=1),$ (5)
in the limit $R\propto\sqrt{N}\uparrow\infty$. Here, the radius $R$ of the fuzzy sphere is interpreted as the radius of compactification in the radial quantization of a CFT with $\Delta_{\mathrm{o}}$ the scaling dimension of the CFT operator labeled by $\mathrm{o}$ and $c$ the system dependent speed of ‘‘light’’ in the CFT.
If we switch on a small perturbation $\delta h=h-h_{\mathrm{c}}\neq 0$ in $\widehat{H}_{\mathrm{fzs}}(R)$ holding $\delta u=\delta v=0$ fixed, the energy shift to first-order in perturbation theory is captured by adding the single relevant scalar operator $\varepsilon$ (see Table 1) with the coupling $g_{\varepsilon}(h)$ together with irrelevant operators to the (2+1)D O(3) Wilson-Fisher critical point. Motivated by first-order conformal perturbation theory (CPT) Lao and Rychkov (2023); Läuchli et al. (2025), we make the scaling Ansatz
$\delta E_{\mathrm{o}}(R,h)=\frac{c}{R}\,\Delta_{\mathrm{o}}(R)+g_{\varepsilon}(h)\,f_{\mathrm{o}\varepsilon\mathrm{o}}(R).$ (6)
Here $\Delta_{\mathrm{o}}(R)=\Delta_{\mathrm{o}}+\mathcal{O}(R^{-\omega})$ deviate from its value at quantum criticality due to the presence of the leading irrelevant operator perturbing the CFT with the scaling exponent $\omega>0$. The value of $h_{\mathrm{c}}$ at the quantum critical point is identified by extrapolating the value of the root to $g_{\varepsilon}(h)=0$ to its limiting value when $R\uparrow\infty$. The interpretation of $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)=f_{\mathrm{o}\varepsilon\mathrm{o}}+\mathcal{O}(R^{-\omega})$ in the limit $R\uparrow\infty$ is the following. (i) It is the (universal) operator product expansion coefficient corresponding to fusing operators $\mathrm{o}$ and $\varepsilon$ into $\mathrm{o}$, if $\mathrm{o}$ labels a scalar primary field. (ii) It is proportional to $f_{\mathrm{o}_{\mathrm{p}}\varepsilon\mathrm{o}_{\mathrm{p}}}(R)$ with the propotionality constant a known function of $\Delta_{\mathrm{o}_{\mathrm{p}}}$ and $\Delta_{\varepsilon}$ if $\mathrm{o}$ labels the descendant of a primary field labeled by $\mathrm{o}_{\mathrm{p}}$ Lao and Rychkov (2023); Läuchli et al. (2025). (iii) Finally, it is a combination of operator-product expansion coefficients Läuchli et al. (2025) if $\mathrm{o}$ labels a field with non-vanishing conformal spin.
Our strategy to establish the existence of the (2+1)D O(3) Wilson-Fisher critical point from exact diagonalization of $\widehat{H}_{\mathrm{fzs}}(R)$ proceeds in two steps. First, we choose two eigenstates (CFT operators) of $\widehat{H}_{\mathrm{fzs}}(R)$ labeled by $\mathrm{o}_{1}$ and $\mathrm{o}_{2}$, for which estimates of $\Delta_{\mathrm{o}_{i}}$ and $f_{\mathrm{o}_{i}\varepsilon\mathrm{o}_{i}}$ with $i=1,2$ at the (2+1)D O(3) Wilson-Fisher critical point are known, to solve for $c$ and $g_{\varepsilon}(h)$ from the pair of equations (6) they generate. Second, for any eigenstate (CFT operator) of $\widehat{H}_{\mathrm{fzs}}(R)$ labeled by $\mathrm{o}$, we extract $\Delta_{\mathrm{o}}(R)$ by using the extracted value of $c$ in Eq. (6) with $g_{\varepsilon}(h)=0$ and then extract $f_{\mathrm{o}\varepsilon\mathrm{o}}$ from the dependence on $R$ of Eq. (6) with $g_{\varepsilon}(h)\neq 0$.
Results for $c$ and $h_{\mathrm{c}}$.— To determine $c$ and $g_{\varepsilon}(h)$, we choose the pair of eigenstates labeled by $\mathrm{o}_{1}\equiv\sigma$ and $\mathrm{o}_{2}\equiv\partial_{\mu}\sigma$, where $\sigma$ labels the lowest eigenenergy of $\widehat{H}_{\mathrm{fzs}}(R)$ with the quantum numbers $S=1^{-}$ and $L=0$, while $\partial_{\mu}\sigma$ labels the lowest eigenenergy of $\widehat{H}_{\mathrm{fzs}}(R)$ with the quantum numbers $S=1^{-}$, $L=1$. In CFT jargon, $\sigma$ labels the most relevant primary field and $\partial_{\mu}\sigma$ is its first descendent. We insert in Eq. (6) $\mathrm{o}_{1}\equiv\sigma$ and $\mathrm{o}_{2}\equiv\partial_{\mu}\sigma$, the conformal bootstrap (CB) estimates $\Delta_{\sigma}=0.518936$ from Table 1 and $\Delta_{\partial_{\mu}\sigma}=1+\Delta_{\sigma}=1.518936$ together with the CB estimates for the OPE coefficient $f_{\sigma\varepsilon\sigma}=0.525$ and $f_{\partial_{\mu}\sigma\varepsilon\partial_{\mu}\sigma}=f_{\sigma\varepsilon\sigma}\,\mathcal{A}_{\sigma,\varepsilon}$ with $\mathcal{A}_{\sigma,\varepsilon}=1+[\Delta_{\varepsilon}(\Delta_{\varepsilon}-3)/(6\Delta_{\sigma})]$ and $\Delta_{\varepsilon}$ given by its CB value quoted in Table 1. The factor $\mathcal{A}_{\sigma,\varepsilon}$ accounts for the fact that the perturbation affects the first descendent differently from the primary Lao and Rychkov (2023); Läuchli et al. (2025). Solving these equations, we obtain the speed of light $c\sim 0.0122$ and the coupling $g_{\varepsilon}(h)$. We then solve for the root of $g_{\varepsilon}(h)=0$ for each value of $N$, to obtain the finite-size critical point as shown in Fig. 1(a). By extrapolating the root of $g_{\varepsilon}(h)=0$ to $N\uparrow\infty$ as shown in Fig. 1(b), we obtain the critical point of the system as $h_{\mathrm{c}}=14.992$.
Рис. 2: Dependence on $R^{-\omega}$ of the dimensionless number $\Delta_{\mathrm{o}}(R)$ defined by the scaling Ansatz (6) with $h$ chosen to be the root of $g_{\varepsilon}(h)=0$ for each value of $R$. Here, the value $\omega\equiv\Delta_{\epsilon^{\prime}}-3\approx 0.7668$ is deduced from the CB estimate for the scaling dimension $\Delta_{\epsilon^{\prime}}$ of the leading irrelevant CFT perturbation (see Table 1). The limiting value $\Delta_{\mathrm{o}}$ of $\Delta_{\mathrm{o}}(R)$ as $R\uparrow\infty$ is interpreted as the scaling dimension of the CFT operator labeled by $\mathrm{o}$. The choice of the eigenstates labeled by $\mathrm{o}$ is organized by $S^{\pm}$ sectors across panels, and by angular momentum $L$ within each panel. Dashed lines indicate conformal bootstrap predictions with the color code green for primary operators and red for descendants. Solid lines represent exact CFT values as applicable to the stress-energy tensor and the conserved Noether current (using the same color code). Operator names, such as $\varepsilon$ and $\varepsilon_{\mu}$, are indicated in red boxes. Only selected sectors are shown here; the complete spectra are available in the Supplementary Material Sup . Panel (a) displays the $S=0^{+}$ sector, panel (b) the $S=1^{+}$ sector, panel (c) the $S=1^{-}$ sector, and panel (d) the $S=2^{+}$ sector. Note that, except for the $\sigma$ and $\partial_{\mu}\sigma$ levels used to determine $c$ and $h_{\mathrm{c}}$, all the plotted values of $\Delta_{\mathrm{o}}(R)$ as $R\uparrow\infty$ are parameter‑free outputs.
Results for scaling dimensions.— The dependence on $R^{3-\Delta_{\varepsilon^{\prime}}}$ of the dimensionless number $\Delta_{\mathrm{o}}(R)$ defined by the scaling Ansatz (6) with $h$ chosen to be the root of $g_{\varepsilon}(h)=0$ for each value of $R$ is shown in Fig. 2 for a selection of eigenstates of $\widehat{H}_{\mathrm{fzs}}(R)$ obtained from ED and DMRG. Here, $\Delta_{\varepsilon^{\prime}}$ is the scaling dimension of the leading irrelevant perturbation, which is estimated to be $3.7668$ by CB. Our estimates for the scaling exponents $\Delta_{\mathrm{o}}$, i.e., the limiting value of $\Delta_{\mathrm{o}}(R)$ as $R\uparrow\infty$, are reported in Table 1 when $\mathrm{o}$ labels a primary operator whose scaling dimensions is known either from CB or exactly, or in Table 2 (Appendix A). A systematic list for the values of the scaling dimensions for all symmetry sectors can be found in the supplementary material Sup .
Figure 2 reveals signatures of emergent conformal symmetry that includes the expected descendant structure, the stress energy tensor, and the conserved SO(3) Noether currents Zhu et al. (2023).
The CB value of the scaling dimension $\Delta_{\epsilon}$ for the primary operator $\varepsilon$ corresponding to the quantum numbers $S=0^{+}$ and $L=0$ is consistent with its ED lower bound and DRMG upper bound from Table 1, respectively. The exact value of the scaling dimension $\Delta_{T_{\mu\nu}}=3$ for the stress-energy tensor $T_{\mu\nu}$ corresponding to the quantum numbers $S=0^{+}$ and $L=2$ is consistent with the ED lower bound and the DMRG upper bound from Table 1. The exact value of the scaling dimension $\Delta_{j_{\mu}}=2$ for the conserved SO(3) Noether current $j_{\mu}$ corresponding to the quantum numbers $S=1^{+}$, $L=1$, that reflects the internal SO(3) symmetry is also consistent with the ED and DMRG estimates from Table 1. Hereto, the CB predictions for the scaling dimensions of the rank-2 $t_{(2)}$ and rank-4 $t_{(4)}$ O(3) primaries are consistent with the ED and DMRG estimates from Table 1. Finally, according to Fig. 2 and Table 2 (Appendix A), there is no CB prediction for the scaling dimension $\Delta_{\sigma_{(3)}}\approx 2.124$ where $\sigma_{(3)}$ corresponds to the quantum numbers $S=3^{-}$ [rank 3 O(3) pseudotensor] and $L=0$ [SO(3) scalar].
In addition, a variety of irrelevant primaries appear, including dangerously irrelevant Amit and Peliti (1982) ones. Among these, the most significant are $\varepsilon^{\prime}$, a scalar under both O(3) and SO(3) (see Appendix B and supplementary material Sup ), i.e., a Laurent expansion in powers of $1/S$ of the scaling dimensions of some low-lying primaries in the internal SO(3) symmetry sector labeled by the quantum number $S\gg 1$. Our data agree with this large $S$ expansion, in particular for the low-$L$ primaries. They are consistent with the existence of both a gapless, linearly dispersing phonon mode, and of a dispersive gapped Goldstone mode.
Рис. 3: Dependence on $R^{-\omega}$ with $\omega\equiv\Delta_{\epsilon^{\prime}}-3\approx 0.7668$ of the coefficient $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)=f_{\mathrm{o}\varepsilon\mathrm{o}}+\mathcal{O}(R^{-\omega})$ defined by Eq. (7) with $\mathrm{o}$ chosen in the $S=0^{\pm}$ symmetry sectors. Solid lines denote CB estimates. Dashed red lines include CPT-derived descendant factors applied to the bootstrap data Lao and Rychkov (2023). The states (CFT operators) selected by $\mathrm{o}$ include $\varepsilon$, $\varepsilon_{\mu}$, $T_{\mu\nu}$.
Results for OPE coefficients.— We extract $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)$ in the scaling Ansatz (6) with the help of
$f_{\mathrm{o}\varepsilon\mathrm{o}}(R)=\left.\frac{\partial\,\delta\,E_{\mathrm{o}}(R)}{\partial\,g_{\varepsilon}(R)}\right|^{\,}_{g_{\varepsilon}(R)=0}.$ (7)
In Fig. 3, we plot the dependence on $R^{-\omega}$ of $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)$ for some of these OPE coefficients and we compare their limiting values as $R\uparrow\infty$ to the available CB data. Other OPE coefficients are plotted and tabulated in the Supplementary Material.
Discussion and outlook.— We have traded a square lattice hosting on each one of its $N$ sites a quantum rotor with a quantum Hamiltonian exhibiting a quantum critical point in the (2+1)-dimensional O(3) Wilson-Fisher universality class for Hamiltonian (4) by placing interacting fermions with $(3+1)$ flavors on a fuzzy sphere. Our method remains computationally tractable using ED and DMRG even for a maximum of $N=26$ interacting fermions and provides evidences for an emergent conformal symmetry consistent with the (2+1)-dimensional O(3) Wilson-Fisher universality class. We identify 24 primary operators labeled by the index $\mathrm{o}$ in Table 1 and Table 2 (Appendix A) and compute their scaling dimensions $\Delta_{\mathrm{o}}$ as well as their diagonal OPE coefficients $f_{\mathrm{o}\varepsilon\mathrm{o}}$. The values that we find agrees with published ones and confirm predictions from the large-$S$ expansion Monin et al. (2017); Cuomo (2020); Cuomo et al. (2021) (see Appendix B).
While analyzing the (2+1)-dimensional O(3) Wilson-Fisher CFT, we identified a weakly irrelevant operator $\varepsilon_{\mu}$ relevant to the criticality of staggered dimerized antiferromagnets. This operator transforms as a pseudoscalar under O(3) and a pseudovector under Lorentz [SO(3)] symmetry, originating from a cubic two-derivative term in the free theory, reminiscent of the topological $\theta$-term. Its behavior in the presence of a boundary remains an intriguing question.
Our framework naturally generalizes to the study of $\mathrm{O}(N)$ Wilson-Fisher quantum criticality, with explicit Hamiltonian constructions provided in the supplementary material Sup . Furthermore, it enables systematic exploration of models exhibiting richer internal symmetries, such as $\mathrm{O}(n_{1})\times\mathrm{O}(n_{2})$ Nelson et al. (1974). This versatility grants access to both Wilson-Fisher and biconal fixed points within a unified approach. By introducing anisotropic terms that break $\mathrm{O}(n)$ symmetry down to a discrete hypercubic subgroup, our method also allows investigation of the stable cubic universality class for $n\geq 3$.
Благодарности.
Acknowledgments.— The authors thank F Mila, Slava Rychkov and Gabriele Cuomo for useful discussions. LH acknowledges the Tremplin funding from CNRS Physique. AD acknowledges useful discussions with João Penedones, Johan Henriksson and Giuseppe Mussardo. AD thanks IHES, Paris for hospitality, where these results were first presented by AD at the workshop "Fuzzy Sphere Meets Conformal Bootstrap June 2-6, 2025.
Note added.— While finishing this article, we became aware of an independent related work Guo et al. .
Приложение A Primary operators
Table 2 presents the primary operators whose scaling dimensions we have computed, for which conformal bootstrap results are not currently available.
Таблица 2: Selected primary operators with the same conventions as in Table 1.
$S$ $L$ II $\mathrm{o}$ $\Delta$(ED) $\Delta$(DMRG)
$0^{-}$ 1 1 $\varepsilon_{\mu}$ 2.96067 (12) 3.2396 (18)
$1^{+}$ 1 3 $\phi_{\mu}$ 3.70828 (12) 3.75703 (18)
$1^{-}$ 0 3 $\sigma^{\prime}$ 3.87329 (12) -
$1^{-}$ 1 2 $\sigma_{\mu}$ 2.87055 (12) 3.04373 (26)
$2^{+}$ 0 3 $t_{(2)}^{\prime}$ 3.67128 (12) 3.68121 (16)
$2^{+}$ 0 4 $t_{(2)}^{\prime\prime}$ 4.31261 (12) -
$2^{+}$ 2 1 $t^{\prime}_{(2)\mu\nu}$ 3.02538 (12) 3.02825 (26)
$2^{-}$ 1 1 $t_{(2)\mu}$ 2.68970 (12) 2.78187 (26)
$2^{-}$ 2 1 $t^{\prime}_{(2)\mu\nu}$ 3.29489 (12) 3.5341 (24)
$3^{+}$ 1 1 $\chi_{(3)\mu}$ 3.68769 (12) 3.75082 (26)
$3^{+}$ 2 1 $\chi_{(3)\mu\nu}$ 4.22326 (12) -
$3^{-}$ 0 1 $\sigma_{(3)}$ 2.17310 (12) 2.12486 (26)
$3^{-}$ 0 3 4.88895 (11) -
$3^{-}$ 2 1 $\sigma_{(3)\mu\nu}$ 4.02221 (12) 3.9825 (24)
$4^{+}$ 0 3 6.25369 (11) -
$4^{+}$ 2 1 $t_{(4)\mu\nu}$ 5.17757 (11) -
$4^{-}$ 1 1 4.85139 (11) 4.87182 (26)
$4^{-}$ 2 1 5.29724 (11) -
[…]
Приложение B Large-$S$ expansion
The large-$S$ expansion provides a controlled semiclassical framework for computing the scaling dimensions of local operators in a CFT with internal O(3) symmetry when these local operators transform according to the $2S+1$-dimensional irreducible representation of O(3). In the sector of the CFT with $S\gg 1$, some scaling dimensions can be associated with the dispersions of two weakly interacting fields associated to the explicit symmetry breaking of the internal symmetry SO(3) down to SO(2). One is massless and called a phonon mode. The other is gapful and called a gapped Goldstone mode.
Three key predictions from the large-$S$ expansion are the following when spacetime is three dimensional.
1. Lowest scalar primaries. The scaling dimension of the lowest scalar $(L=0$) operator in each $S\gg 1$ sector obeys the semiclassical expansion
$\begin{split}\Delta_{0}(S^{\kappa(S)},L=0)=&\,\alpha\,S^{3/2}+\beta\,S^{1/2}-0.0937256+\gamma\,S^{-1/2}\\ &\,+\mathcal{O}(1/S),\end{split}$ (8)
where $\kappa(S)=(-1)^{S}$ determines the parity. The Wilson coefficients $\alpha=0.31076$, $\beta=0.29818$, and $\gamma=0.00370$ are obtained by fitting the lowest level in the sectors $1^{-}$, $2^{+}$, $4^{+}$ to bootstrap data for $\sigma$, $t_{2}$, and $t_{4}$.
2. Gapped Goldstone modes. The scaling dimensions of the state with spin $S\gg 1$, parity $\kappa(S+1)=-\kappa(S)$, and non-vanishing angular momentum $L\geq 1$ obeys the semiclassical expansion
$\begin{split}\Delta(S^{\kappa(S+1)},L)=&\,\Delta_{0}(S^{\kappa(S)},L=0)+\mu(S)+\chi\,\frac{L(L+1)}{2\mu(S)}\\ &\,+\mathcal{O}\left(\frac{L^{4}}{\mu^{3}}\right),\end{split}$ (9a)
where
$\mu(S):=\frac{\partial\Delta_{0}(S,L=0)}{\partial S}$ (9b)
is the chemical potential that breaks explicitly the internal continuous symmetry from SO(3) to SO(2). The term $L(L+1)/(2\mu)$ encodes the dispersion of the gapped Goldstone boson. The value of the Wilson coefficient $\chi$ is obtained by fitting the lowest scaling dimension in the sector $1^{+}$ to the value 2 of the scaling dimension of the Noether current $j_{\mu}$.
3. Gapless phonon modes. The scaling dimension of the state with the quantum numbers $S\gg 1$ for the spin, parity $\kappa(S)=(-1)^{S}$, and the angular momentum $L=0,1,\cdots$ obeys the semiclassical expansion
$\Delta_{0}(S^{\kappa(S)},L)=\Delta_{0}(S^{\kappa(S)},L=0)+\sum^{\,}_{L}n^{\,}_{L}\omega^{\,}_{L},$ (10a)
where
$\omega^{\,}_{L}=\sqrt{L(L+1)/2},$ (10b)
is the energy of the phonon with angular momentum $L$ and the integer $n^{\,}_{L}$ is its occupancy. The value $L=1$ correspond to a phonon descendant, while the values $L\geq 2$ correspond to phonon primaries.
The agreement between these theoretical predictions and fuzzy-sphere numerical results (shown in Fig. 4) provides strong evidence for the presence of both gapped Goldstone modes and gapless phonon modes in the (2+1)-dimensional O(3) Wilson-Fisher CFT, validating the large-$S$ expansion framework.
Список литературы

### v2

@@
 Conformal Data for the O(3) Wilson-Fisher CFT from Fuzzy Sphere Realization of Quantum Rotor Model
-Arjun Dey Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Loic Herviou Univ.
-Grenoble Alpes, CNRS, LPMMC, 38000 Grenoble, France    Christopher Mudry Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Andreas Martin Läuchli Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polyetchnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland
+Arjun Dey Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polytechnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Loic Herviou Univ.
+Grenoble Alpes, CNRS, LPMMC, 38000 Grenoble, France    Christopher Mudry Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polytechnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland    Andreas Martin Läuchli Laboratory for Theoretical and Computational Physics, PSI Center for Scientific Computing, Theory and Data, 5232 Villigen PSI, Switzerland Institute of Physics, École Polytechnique Fédérale de Lausane (EPFL), 1015 Lausane, Switzerland
 Аннотация
-We present a model for strongly interacting fermions with internal O(3) symmetry on the fuzzy-sphere that (i) preserves the rotational symmetry of the fuzzy sphere and (ii) undergoes a quantum phase transition in the (2+1)-dimensional O(3) Wilson–Fisher universality class.
+We present a model for strongly interacting fermions with internal O(3) symmetry on the fuzzy sphere that (i) preserves the rotational symmetry of the fuzzy sphere and (ii) undergoes a quantum phase transition in the (2+1)-dimensional O(3) Wilson-Fisher universality class.
 Using exact diagonalization (ED) and density matrix renormalization group (DMRG), we locate the quantum critical point via conformal perturbation theory and obtain scaling dimensions from finite-size spectra.
@@
 (2025) provides a direct Hamiltonian approach to studying (2+1)D CFTs through diagonalization of quantum systems on a sphere, offering access to both spectra and correlation functions while preserving full symmetries.
-A prior fuzzy-sphere study of the (2+1)D CFT with internal O(3) symmetry confirmed feasibility but was limited by Hilbert-space growth and identified only a limited number of operators Han et al.
+A prior fuzzy sphere study of the (2+1)D CFT with internal O(3) symmetry confirmed feasibility but was limited by Hilbert-space growth and identified only a limited number of operators Han et al.
 (2024).
-Here, we introduce a fuzzy-sphere realization (FZR) of a truncated quantum rotor model (TQRM) Sachdev (2011); Zhu et al.
-(2023) that retains the full internal O(3) and spatial SO(3) symmetry of the (2+1)-dimensional O(3) Wilson–Fisher universality class and is tractable to ED and DMRG.
+Here, we introduce a fuzzy sphere realization (FZR) of a truncated quantum rotor model (TQRM) Sachdev (2011); Zhu et al.
+(2023) that retains the full internal O(3) and spatial SO(3) symmetry of the (2+1)-dimensional O(3) Wilson-Fisher universality class and is tractable to ED and DMRG.
 Building on the well‐known correspondence between O(3) quantum rotors and O(3) quantum criticality Sachdev (2011), the idea is to use a conformal map to identify a compactification of two-dimensional space with the unit 2-sphere $\mathbb{S}^{2}$ and to leverage the state-operator correspondence to extract conformal data.
@@
 Quantum rotor model.— A rigid rotor, a particle of mass $\mu$ constrained to move on the surface of a sphere of radius $R$, serves as the prototypical quantum model of rotational motion.
-Its energy is purely kinetic, arising from quantized angular momentum with moment of inertia $I=\mu\,R^{2}>0$.
+Its energy is purely kinetic, arising from quantized angular momenta with moment of inertia $I=\mu\,R^{2}>0$.
 On a square lattice, an array of such rotors with nearest-neighbor ferromagnetic Heisenberg coupling $J>0$ realizes an O(3) magnetic transition when the dimensionless coupling IJI\,J is tuned to quantum criticality Sachdev (2011).
@@
 […]
-Fuzzy-sphere many-body electronic Hamiltonian.— To access conformal data, we map the quantum rotor model to a system of $N$ fermions with internal degrees of freedom, moving on a sphere Haldane (1983) and projected to the lowest Landau level (LLL).
-In this limit, the noncommuting coordinate operators generate the fuzzy-sphere algebra Madore (1992), thereby providing a natural short-distance cutoff (the magnetic length) that preserves rotational symmetry.
-We focus on the case where each fermion has $F=4$ flavors and the system reside at quarter filling.
+Fuzzy sphere many-body electronic Hamiltonian.— To access conformal data, we map the quantum rotor model to a system of $N$ fermions with internal degrees of freedom, moving on a sphere Haldane (1983) and projected to the lowest Landau level (LLL).
+In this limit, the noncommuting coordinate operators generate the fuzzy sphere algebra Madore (1992), thereby providing a natural short-distance cutoff (the magnetic length) that preserves rotational symmetry.
+We focus on the case where each fermion has $F=4$ flavors and the system resides at quarter filling.
 At quarter filling, strong interactions enforce (on average) single occupancy per LLL orbital, analogous to the large-$U$ limit of the Hubbard model.
-The microscopic fuzzy-sphere Hamiltonian that we shall study numerically is defined to be
+The microscopic fuzzy sphere Hamiltonian that we shall study numerically is defined to be
 $\widehat{H}_{\mathrm{fzs}}:=\widehat{P}_{\mathrm{LLL}}\left(u\,\widehat{H}_{\mathrm{Hub}}-v\,\widehat{H}_{\mathrm{Heis}}+h\,\widehat{H}_{\mathrm{trans}}\right)\widehat{P}_{\mathrm{LLL}}.$ (4)
@@
 Таблица 1: Summary of selected primary operators.
-$S$ labels the internal SO(3) spin representation, $L$ the SO(3) angular momentum on the fuzzy sphere, and $\pm$ the $\mathbb{Z}_{2}$ parity (improper part) of the internal O(3) symmetry, so that states are labeled by $S^{\pm}$ and $L$.
+$S$ labels the internal SO(3) spin representation, $L$ the SO(3) angular momenta on the fuzzy sphere, and $\pm$ the $\mathbb{Z}_{2}$ parity (improper part) of the internal O(3) symmetry, so that states are labeled by $S^{\pm}$ and $L$.
 The index I orders energy eigenstates within each symmetry sector.
@@
 (2002).
-Symmetries.— Hamiltonian (4) exhibits three symmetries: (a) spatial rotational symmetry SO(3), arising from conservation of angular momentum with the quantum number $L$; (b) internal SO(3) spin symmetry with the quantum number $S$; (c) and a $\mathbb{Z}_{2}$ symmetry corresponding to the conservation of the parity of the number of electrons occupying $p$ orbitals in internal spin space.
+Symmetries.— Hamiltonian (4) exhibits three symmetries: (a) spatial rotational symmetry SO(3), arising from conservation of angular momenta with the quantum number $L$; (b) internal SO(3) spin symmetry with the quantum number $S$; (c) and a $\mathbb{Z}_{2}$ symmetry corresponding to the conservation of the parity of the number of electrons occupying $p$ orbitals in internal spin space.
 The quantum number for the $\mathrm{SO}(3)\times\mathbb{Z}_{2}\cong\mathrm{O}(3)$ is denoted by $S^{\pm}$, where the superscript denotes the $\mathbb{Z}_{2}$ parity.
@@
 in the limit $R\propto\sqrt{N}\uparrow\infty$.
-Here, the radius $R$ of the fuzzy sphere is interpreted as the radius of compactification in the radial quantization of a CFT with $\Delta_{\mathrm{o}}$ the scaling dimension of the CFT operator labeled by $\mathrm{o}$ and $c$ the system dependent speed of ‘‘light’’ in the CFT.
-If we switch on a small perturbation $\delta h=h-h_{\mathrm{c}}\neq 0$ in $\widehat{H}_{\mathrm{fzs}}(R)$ holding $\delta u=\delta v=0$ fixed, the energy shift to first-order in perturbation theory is captured by adding the single relevant scalar operator $\varepsilon$ (see Table 1) with the coupling $g_{\varepsilon}(h)$ together with irrelevant operators to the (2+1)D O(3) Wilson-Fisher critical point.
+Here, the radius $R$ of the fuzzy sphere is interpreted as the radius of compactification in the radial quantization of a CFT, $\Delta_{\mathrm{o}}$ the scaling dimension of the CFT operator labeled by $\mathrm{o}$ and $c$ the system dependent speed of ‘‘light’’ in the CFT.
+If we switch on a small perturbation $\delta h=h-h_{\mathrm{c}}\neq 0$ in $\widehat{H}_{\mathrm{fzs}}(R)$ holding $\delta u=\delta v=0$ fixed, the energy shift to first-order in perturbation theory is captured by adding the single relevant scalar operator $\varepsilon$ (see Table 1) with the coupling $g_{\varepsilon}(h)$, together with irrelevant operators to the (2+1)D O(3) Wilson-Fisher critical point.
 Motivated by first-order conformal perturbation theory (CPT) Lao and Rychkov (2023); Läuchli et al.
@@
 2: Dependence on $R^{-\omega}$ of the dimensionless number $\Delta_{\mathrm{o}}(R)$ defined by the scaling Ansatz (6) with $h$ chosen to be the root of $g_{\varepsilon}(h)=0$ for each value of $R$.
-Here, the value $\omega\equiv\Delta_{\epsilon^{\prime}}-3\approx 0.7668$ is deduced from the CB estimate for the scaling dimension $\Delta_{\epsilon^{\prime}}$ of the leading irrelevant CFT perturbation (see Table 1).
+Here, the value $\omega\equiv\Delta_{\varepsilon^{\prime}}-3\approx 0.7668$ is deduced from the CB estimate for the scaling dimension $\Delta_{\varepsilon^{\prime}}$ of the leading irrelevant CFT perturbation (see Table 1).
 The limiting value $\Delta_{\mathrm{o}}$ of $\Delta_{\mathrm{o}}(R)$ as $R\uparrow\infty$ is interpreted as the scaling dimension of the CFT operator labeled by $\mathrm{o}$.
-The choice of the eigenstates labeled by $\mathrm{o}$ is organized by $S^{\pm}$ sectors across panels, and by angular momentum $L$ within each panel.
+The choice of the eigenstates labeled by $\mathrm{o}$ is organized by $S^{\pm}$ sectors across panels; and by angular momenta $L$ within each panel.
 Dashed lines indicate conformal bootstrap predictions with the color code green for primary operators and red for descendants.
@@
 Here, $\Delta_{\varepsilon^{\prime}}$ is the scaling dimension of the leading irrelevant perturbation, which is estimated to be $3.7668$ by CB.
-Our estimates for the scaling exponents $\Delta_{\mathrm{o}}$, i.e., the limiting value of $\Delta_{\mathrm{o}}(R)$ as $R\uparrow\infty$, are reported in Table 1 when $\mathrm{o}$ labels a primary operator whose scaling dimensions is known either from CB or exactly, or in Table 2 (Appendix A).
+Our estimates for the scaling dimensions $\Delta_{\mathrm{o}}$, obtained from the largest available ED and DMRG system sizes rather than from an explicit extrapolation to $R\uparrow\infty$, are reported in Table 1 when $\mathrm{o}$ labels a primary operator whose scaling dimension is known either from the conformal bootstrap or exactly, and in Table 2 (Appendix A) otherwise.
 A systematic list for the values of the scaling dimensions for all symmetry sectors can be found in the supplementary material Sup .
 Figure 2 reveals signatures of emergent conformal symmetry that includes the expected descendant structure, the stress energy tensor, and the conserved SO(3) Noether currents Zhu et al. (2023).
-The CB value of the scaling dimension $\Delta_{\epsilon}$ for the primary operator $\varepsilon$ corresponding to the quantum numbers $S=0^{+}$ and $L=0$ is consistent with its ED lower bound and DRMG upper bound from Table 1, respectively.
+The CB value of the scaling dimension $\Delta_{\varepsilon}$ for the primary operator $\varepsilon$ corresponding to the quantum numbers $S=0^{+}$ and $L=0$ is consistent with its ED lower bound and DMRG upper bound from Table 1, respectively.
 The exact value of the scaling dimension $\Delta_{T_{\mu\nu}}=3$ for the stress-energy tensor $T_{\mu\nu}$ corresponding to the quantum numbers $S=0^{+}$ and $L=2$ is consistent with the ED lower bound and the DMRG upper bound from Table 1.
 The exact value of the scaling dimension $\Delta_{j_{\mu}}=2$ for the conserved SO(3) Noether current $j_{\mu}$ corresponding to the quantum numbers $S=1^{+}$, $L=1$, that reflects the internal SO(3) symmetry is also consistent with the ED and DMRG estimates from Table 1.
-Hereto, the CB predictions for the scaling dimensions of the rank-2 $t_{(2)}$ and rank-4 $t_{(4)}$ O(3) primaries are consistent with the ED and DMRG estimates from Table 1.
+In addition, the CB predictions for the scaling dimensions of the rank-2 $t_{(2)}$ and rank-4 $t_{(4)}$ O(3) primaries are consistent with the ED and DMRG estimates from Table 1.
 Finally, according to Fig.
@@
 In addition, a variety of irrelevant primaries appear, including dangerously irrelevant Amit and Peliti (1982) ones.
-Among these, the most significant are $\varepsilon^{\prime}$, a scalar under both O(3) and SO(3) (see Appendix B and supplementary material Sup ), i.e., a Laurent expansion in powers of $1/S$ of the scaling dimensions of some low-lying primaries in the internal SO(3) symmetry sector labeled by the quantum number $S\gg 1$.
+Among these, the most significant are $\varepsilon^{\prime}$, a scalar under both O(3) and SO(3) [see Fig.
+2(a)], and $\varepsilon_{\mu}$ a pseudoscalar under O(3) and a pseudovector under SO(3) [see Fig.
+2(b)].
+These operators govern leading corrections to scaling in lattice simulations Fritz et al.
+(2011); Ma et al.
+(2018).
+Notably, $\varepsilon_{\mu}$ provides insight into a long-standing discrepancy in the critical exponents of columnar versus staggered dimerized antiferromagnets Fritz et al.
+(2011); Kao et al.
+(2012); Yasuda and Todo (2013); Ma et al.
+(2018), once misattributed to different universality classes Wenzel et al.
+(2008).
+Our finite-size extrapolation using a single power law with exponent $0.7668$ yields the scaling dimension $\Delta\approx 3.931$, in reasonable agreement with the five-loop $\epsilon$-expansion result of 3.78796 Henriksson et al.
+(2025).
+The identification of this operator as a three-body term with two derivatives and dimension 3.5 in the free O(3) theory further corroborates predictions in the literature Fritz et al.
+(2011).
+At last, the spectrum of scaling exponent that we found numerically is compared with those obtained from the large-$S$ expansion [see Refs.
+Cuomo et al.
+(2021); Cuomo (2020); Sup and Appendix B], i.e., a Laurent expansion in powers of $1/S$ of the scaling dimensions of some low-lying primaries in the internal SO(3) symmetry sector labeled by the quantum number $S\gg 1$.
 Our data agree with this large $S$ expansion, in particular for the low-$L$ primaries.
-They are consistent with the existence of both a gapless, linearly dispersing phonon mode, and of a dispersive gapped Goldstone mode.
+They are consistent with the existence of both a gapless, linearly dispersing phonon mode and of a dispersive gapped Goldstone mode.
 Рис.
-3: Dependence on $R^{-\omega}$ with $\omega\equiv\Delta_{\epsilon^{\prime}}-3\approx 0.7668$ of the coefficient $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)=f_{\mathrm{o}\varepsilon\mathrm{o}}+\mathcal{O}(R^{-\omega})$ defined by Eq.
+3: Dependence on $R^{-\omega}$ with $\omega\equiv\Delta_{\varepsilon^{\prime}}-3\approx 0.7668$ of the coefficient $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)=f_{\mathrm{o}\varepsilon\mathrm{o}}+\mathcal{O}(R^{-\omega})$ defined by Eq.
 (7) with $\mathrm{o}$ chosen in the $S=0^{\pm}$ symmetry sectors.
@@
 3, we plot the dependence on $R^{-\omega}$ of $f_{\mathrm{o}\varepsilon\mathrm{o}}(R)$ for some of these OPE coefficients and we compare their limiting values as $R\uparrow\infty$ to the available CB data.
-Other OPE coefficients are plotted and tabulated in the Supplementary Material.
-Discussion and outlook.— We have traded a square lattice hosting on each one of its $N$ sites a quantum rotor with a quantum Hamiltonian exhibiting a quantum critical point in the (2+1)-dimensional O(3) Wilson-Fisher universality class for Hamiltonian (4) by placing interacting fermions with $(3+1)$ flavors on a fuzzy sphere.
+Other OPE coefficients are plotted and tabulated in the Supplementary Material Sup .
+Discussion and outlook.— We have mapped the square-lattice quantum rotor model, whose Hamiltonian exhibits a quantum critical point in the (2+1)-dimensional O(3) Wilson-Fisher universality class, onto the fuzzy sphere Hamiltonian (4) by placing interacting fermions with four internal flavors on the sphere.
 Our method remains computationally tractable using ED and DMRG even for a maximum of $N=26$ interacting fermions and provides evidences for an emergent conformal symmetry consistent with the (2+1)-dimensional O(3) Wilson-Fisher universality class.
@@
 where $\kappa(S)=(-1)^{S}$ determines the parity.
-The Wilson coefficients $\alpha=0.31076$, $\beta=0.29818$, and $\gamma=0.00370$ are obtained by fitting the lowest level in the sectors $1^{-}$, $2^{+}$, $4^{+}$ to bootstrap data for $\sigma$, $t_{2}$, and $t_{4}$.
-2. Gapped Goldstone modes. The scaling dimensions of the state with spin $S\gg 1$, parity $\kappa(S+1)=-\kappa(S)$, and non-vanishing angular momentum $L\geq 1$ obeys the semiclassical expansion
+The Wilson coefficients $\alpha=0.31076$, $\beta=0.29818$, and $\gamma=0.00370$ are obtained by fitting the lowest level in the sectors $1^{-}$, $2^{+}$, $4^{+}$ to bootstrap data for $\sigma$, $t_{(2)}$, and $t_{(4)}$.
+2. Gapped Goldstone modes. The scaling dimensions of the state with spin $S\gg 1$, parity $\kappa(S+1)=-\kappa(S)$, and non-vanishing angular momenta $L\geq 1$ obeys the semiclassical expansion
 $\begin{split}\Delta(S^{\kappa(S+1)},L)=&\,\Delta_{0}(S^{\kappa(S)},L=0)+\mu(S)+\chi\,\frac{L(L+1)}{2\mu(S)}\\ &\,+\mathcal{O}\left(\frac{L^{4}}{\mu^{3}}\right),\end{split}$ (9a)
@@
 Gapless phonon modes.
-The scaling dimension of the state with the quantum numbers $S\gg 1$ for the spin, parity $\kappa(S)=(-1)^{S}$, and the angular momentum $L=0,1,\cdots$ obeys the semiclassical expansion
+The scaling dimension of the state with the quantum numbers $S\gg 1$ for the spin, parity $\kappa(S)=(-1)^{S}$, and the angular momenta $L=0,1,\cdots$ obeys the semiclassical expansion
 $\Delta_{0}(S^{\kappa(S)},L)=\Delta_{0}(S^{\kappa(S)},L=0)+\sum^{\,}_{L}n^{\,}_{L}\omega^{\,}_{L},$ (10a)
@@
 $\omega^{\,}_{L}=\sqrt{L(L+1)/2},$ (10b)
-is the energy of the phonon with angular momentum $L$ and the integer $n^{\,}_{L}$ is its occupancy.
+is the energy of the phonon with angular momenta $L$ and the integer $n^{\,}_{L}$ is its occupancy.
 The value $L=1$ correspond to a phonon descendant, while the values $L\geq 2$ correspond to phonon primaries.
-The agreement between these theoretical predictions and fuzzy-sphere numerical results (shown in Fig.
+The agreement between these theoretical predictions and fuzzy sphere numerical results (shown in Fig.
 4) provides strong evidence for the presence of both gapped Goldstone modes and gapless phonon modes in the (2+1)-dimensional O(3) Wilson-Fisher CFT, validating the large-$S$ expansion framework.